OpenStudy (anonymous):

4sqrt{2}-6sqrt{8} __how do i simplify this radical expression

OpenStudy (anonymous):

same as before noting that \[8=4\times 2\] so \[\sqrt{8}=2\sqrt{2}\]

OpenStudy (anonymous):

that answer s wrong

OpenStudy (anonymous):

4sqrt{2}-6sqrt{8}=4sqrt{2}-6sqrt{2x4}= 4sqrt{2}-6sqrt{2}xsqrt{4}= 4sqrt{2}-6xsqrt{2}x2= 4sqrt{2}-12sqrt{2}=-8sqrt{2}